The rotor of an electric motor has rotational inertia I 4.41 x 103 kg m2 about its central axis. The motor is used to change the crientation of the space probe in which it is mounted. The motor axis is mounted along the central axis of the probe; the probe has rotational inertia L 7.77 kg.m2 about this axis. Calculate the number of revolutions of the rotor required to tum the probe through 44.8 about its central axis. Units Numbe the tolerance is +/-2% Click if you would like to show Work for this question Open Show Work
Answer #1

4. Чх10-3 (360xn)= 7.17 × (43.6 n= 213.81 九etations